OpenStudy (anonymous):

Solve for x using logarithms. Give an exact answer. Do not give a decimal approximation. 53 = 350e^−0.4x

OpenStudy (jdoe0001):

\(\bf \large 53 = 350e^{−0.4x} \quad ?\)

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

\(\bf 53 = 350e^{−0.4x} \implies \cfrac{50}{350}= e^{−0.4x}\\ \textit{recall the log cancellation rule of } log_aa^x = x\\ log_e\left(\cfrac{50}{350}\right) = log_e\left(e^{−0.4x}\right)\) what would that equate to?

OpenStudy (jdoe0001):

notice that \(\bf \textit{log cancellation rule of } log_\color{red}{a}\color{red}{a}^x = x\\ log_e\left(\cfrac{50}{350}\right) = log_\color{red}{e}\left(\color{red}{e}^{−0.4x}\right)\)

jhonyy9 (jhonyy9):

@jdoe0001 ATTENTION !!! the first term is 53 and not 50 !!!